Gas systems weekly project meeting

→ Europe/Zurich
20/1-004 (CERN)

20/1-004

CERN

15
Show room on map
gas systems project meeting
Zoom Meeting ID
62095888225
Host
Beatrice Mandelli
Useful links
Join via phone
Zoom URL

Gas systems 2026-10-01

Key Outcomes

Multiple gas system issues were identified and partially resolved across ATLAS and CMS RPC systems. The RPC mixer at 904 is now operational after pressure troubleshooting. CLOUD software and regulation bugs remain open issues, with full resolution expected by end of next week. On-call (piquet) coverage was discussed for the Christmas period.

 

Incidents & Root Causes

·       ATLAS RPC humidifier: Water level in reservoir dropped to below 40% (alarm threshold at 40%);

·       ATLAS-TGC: a perforated pneumatic tube on the drain valve, which opened fully and redirected all CO₂ flow through the reservoir (ELOG - Atlas TGC alarme PCV4003 a 100%, fuite pneumatique).

·       Root cause: Cold tubes passing through a wall formed ice, which over time crushed the pneumatic tubing; electrical cables share the same conduit, raising further concerns.

·       Profibus card issues: Recurring problems noted; alarms deliberately not cleared pending investigation (example ELOG - Alice MCH alarme Profibus).

·       CLOUD : A regulation error caused water to enter the chamber at >100°C, generating steam and disrupting flow control; membrane has since been replaced.

·       CLOUD automatic drain malfunction: Outlet valve failed to close at the correct level, sending water into the chamber instead of gas.

 

System Status

·       RPC mixer at 904: Now stable — line pressure raised to 100 mbar (was only 30 mbar with 10 rotameters open); flow at ~20 L/h total. 

·       RPC 904 Nitrogen humidifier: Connection planned for Monday; expected operational Monday afternoon.

·       CMS software (PVSS): Non-functional until late August; heating regulation parameters were incorrect and required manual correction one by one.  Most issues resolved; remaining bugs targeted for end of next week.

·       Flashing lights at S62 control room: Started flashing unexpectedly during dismounting; likely a cut cable — action to remove power and physically dismount when available.

 

Open Issues & Risks

·       CLOUD software bugs: Marc (software contact) unavailable; follow-up session planned for Monday.

·       Purifier trend review: Pressure increase observed across all loop points; unclear if fault originated at purifier. Alarm logic delay may be 5 minutes, not 5 seconds — needs verification for CMS and all systems.

·       CMS humidifier flow: Delivering only 15–20 g/h against a 100 g/h full-scale set point; purifier partially bypassed — ordering a replacement recommended.

·       LHCb UT valve/pipe condition: Item on action list with unclear description; possibly relates to valve not opening or pneumatic pipe condition — needs clarification.

·       LHCb UT flow stop behavior: When system is manually stopped, flow stops; needs check against functional analysis to confirm expected behavior.

·       RPC differential pressure review: Econex inspection not yet scheduled — flagged as at risk of being deprioritized if not formally planned.

 

Action Items

·       Louis-Philippe/Gianluca: Investigate and resolve remaining CLOUD software bugs with Marc — Monday session.

·       Andrea: Notify RPC team to maintain line pressure at several hundred mbar minimum.

·       Andrea: Connect nitrogen humidifier Monday; confirm operational by Monday afternoon.

·       Patrick: Remove power to flashing lights at NA62; dismount when opportunity arises.

·       Gianluca: Verified buffer pressure alarm delay unit (5 seconds vs. 5 minutes) in recipe.-> there is an issue in the recipe

·       Andrea->Patrick: Order replacement CMS humidifier purifier (current unit insufficient at minimum flow).

·       Andrea/Maria Cristina: Schedule Econex/RPC valve inspection

·       Patrick/Stefania/Gianluca: Clarify LHCb UT pipe/valve condition item; check functional analysis for manual stop behavior.

·       NA62 Straw: Fabio available for up to two days; Roberto to confirm based on bottle availability.

·       Roberto: Prepare first batch of Christmas piquet schedule; confirm assignments next week.

 

Christmas Holiday On-Call (PK) Coverage

·       Period: 23 Dec – early January (~14 days)

·       Available personnel: Gianluca (from 23rd), Davide, Louis-Philippe, Andrea, Maria-Cristina, Patrick, Beatrice (avoid last 2–3 days of holidays)

·       ~5–6 people available; assignments to be distributed roughly in alphabetical order; final schedule to be confirmed next week.

·       Last regular piquet ends 22nd (maria Cristina); Gianluca's slot finishes 17th.

Gas systems 2026-09-24 

Action Items

·         Gianluca: Investigate PLC logic for minimum valve opening threshold below 10%. CMS-RPC humidifier

·         Maria Cristina: Review purifier pressure trend; assess feasibility of dedicated purifier high-pressure alarm.

·         Gianluca: Verify buffer pressure alarm delay unit (5 seconds vs. 5 minutes) in recipe.

·         Davide/Roberto / Fabio: Proceed with NA62 Straw chamber disconnection; coordinate access permit with Roberto.

·         Stefania (microcirculation): Order mechanical and electrical material; report estimated arrival at next meeting.

·         Davide/Max/GIanluca: Execute Atlas isobutane mixture change starting ~5 October, allowing 3–5 days for stabilization.

·         Davide/Max/Gianluca: Check historical pump test record to identify which pump had the prior current absorption problem ATL-MDT.

Gas systems 2026 09 17

Action Items

·         Gianluca: Check PLC logic for humidifier mass flow controller minimum opening threshold (with Gianluca, back next week) CMS-RPC.

·         Andrea: Order correct-scale mass flow controller for water injection (~20 g/hr max scale based on observed demand).

·         Gianluca/Maria Cristina/Roberto: Review purifier differential pressure trend to determine if early blockage detection is feasible; consider adding dedicated alarm.

·         Andrea/Maria Cristina: Inspect Econex valve seat from Rack 70 for aging/deformation; assess all seats for LS3 replacement. Not yet dismounted

·         Davide/Max: Continue CO2 pressure-drop tests on Column B (Atlas-RPC) — check if drop rate decreases over repeated tests (saturation indicator). -> stable rate

·         Davide: Perform full pump maintenance on all three RPC pumps (gasket material upgrade, needle valve replacement). -> done maintenance p2, non return completely broken. replaced. pump 2 still not ok after. Plan to replace all non-return

 

Gas systems 2026-09-03

Action Items

  • Patrick/Gianluca: Investigate LHCb-UT,  confirm correct logic and pipe  (or valve or pneumatic pipes?) condition.
  • Andrea :  finalize TID and send link.
  • Florian's mixer: Order required materials — planned for next week – who is ordering? Fabio ongoing Patrick for electrical
  • Andrea: uCL electrical part to put in the workshop planning w42-45 empty but full. now in week 46.

 

 

 

 

 

 

 

 

 

piquet:

 

LHCb pump filters to be changed or something (for LS3)-> TICKET

 

O2+H2O analyser to be checked ->  TICKET

 

ATLAS-RPC rack 66 and 69 pressure alarm during purif 2 change over -> TICKET

 

NA62

glitch of CEDAR 6105, is it the keller? -> TICKET to check

 

ATLAS-TGC putifier Col B leak. loosing 1 bar in 1 day -> TICKET Davide check with Argon or CO2

 

CMS-DT pressure sensor PT1303 went to 0 for 20 minutes -> TICKET to follow Andrea

 

ATLAS-RPC purif -> affecting chamber pressure -> R66 check

12/2 ATL-RPC purif 1 some software issue and sms not indicating the right reason for the alarm -> TICKET

 

IR status, spares, .... --> TICKET

test binary gas analyser... -> TICKET

outcome of binary gas analyser test --> TICKET & buy a 2nd one --> Maria Cristina

DP pressure for pump filter in ALICE and LHCb -> TICKET -> @LPh + @Patrick [GASSYSTEMS-903] DP pressure for pump filter increasing - CERN Central Jira  -> Stefania

chiller fluids to uniform between CMS, LHCb and ATLAS -> TICKET Maria Cristina 

 

check if NV in purif is really useful --> TICKET

 

Calibration bottles O2-H2O status

 

ALICE

70 report from Anité [GASSYSTEMS-70] TPC gas quality - filters to Benoit - CERN Central Jira (I don't see the report, can you please add it?)

several alarms blocked: [GASSYSTEMS-628] ALICE alarms blocked - CERN Central Jira

TOF gas mixture analysis -> [GASSYSTEMS-639] TOF gas mixture analysis - CERN Central Jira & ELOG

TRD membrane is connected ? -> [GASSYSTEMS-717] TRD membrane status - CERN Central Jira

P&ID drawing check who is doing -> [GASSYSTEMS-736] ALICE P&ID drawing check - CERN Central Jira

MID swap of pumps during MD because sometime the system stop -> TICKET

 

 

 

ATLAS

MDT frequency converter (need to be replaced TICKET) ELOG - ATLAS MDT Test convertisseur 4101 avec la pompe 4201 

RPC distribution rack regulation valve (TICKET)

RPC secondary mixer (TICKET), info to be given to ATLAS

RPC purif 3 circuit breakers & meeting -> [GASSYSTEMS-718] RPC purif 3 electrical issue - CERN Central Jira --> meeting when?

ATLAS-TGC flow difference --> [GASSYSTEMS-612] ATLAS-TGC flow difference - CERN Central Jira

 

 

CMS

RPC purif test at workshop, PLC needed -> TICKET

R134a insulation - isol'tut to check and see what to do -> ticket IO [GASSYSTEMS-578] R134a insulation - isol'tut to check and see what to do - CERN Central Jira

How much gas flow we need for the RPC purifier? -->  [GASSYSTEMS-579] CMS-RPC How much gas flow we need for the RPC purifiers? - CERN Central Jira @MariaCristina notes to be added

RPC bypass pump too much closed --> [GASSYSTEMS-604] RPC bypass pump too much closed - CERN Central Jira

RPC with CO2 like ATLAS and effect on R134a recuperation -> [GASSYSTEMS-606] RPC with CO2 like ATLAS and effect on R134a recuperation - CERN Central Jira

R134a Upgrade (hardware and software) to be tested -> [GASSYSTEMS-594] R134a Upgrade (hardware and software) to be tested - CERN Central Jira

DT vaisala reading difference between SG and US -> [GASSYSTEMS-677] CMS DT vaisala reading difference between SG and US - CERN Central Jira

CMS-FGR -> TICKET should we add a module. Do we need to buy something? follow-up

CMS-FGR -> TICKET Maria Cristina già fatto per bypass

CMS-FGR -> TICKET flowmeter for Huber

CMS-RPC IR not stable -> elog Maria Cristina -> TICKET

CMS-ID flushing. refurbishing of old racks material and manpower need to be estimated -> TICKET @Gianluca ticket number for project

R134a huber temperature can be controlled by PLC -> TICKET

 

 

LHCb

TS1-> RICH1 leak search (postponed), SciFi some work from others expected

RICH1 what we learn from 1st fill -> [GASSYSTEMS-608] RICH1 what we learn from 1st fill - CERN Central Jira

summary of the C4F10 bottles --> TICKET or ELOG -> [GASSYSTEMS-660] summary of the C4F10 bottles - CERN Central Jira

alarms blocked: [GASSYSTEMS-631] LHCb alarms blocked - CERN Central Jira

RICH2 filter blocking -> [GASSYSTEMS-731] RICH2 pump filter differential pressure sensor not working - CERN Central Jira

P&ID drawing check who is doing -> [GASSYSTEMS-739] LHCb P&ID drawing check - CERN Central Jira

RICH1 cleaning , when? not before mid-september? -> TICKET 

RICH1 remove membrane rack before end of HL-LHC operation and/or change piping-> TICKET

RICH2 injection through recuperated CF4 -> TICKET

fire detection alarms reported by Pieter, more info? but not on us. Cleaning of SG room -> TICKET 

CODEX-b interlock for flammable gas detection -> Patrick

RICH1 leak repair -> December 9th tbc -> TICKET

 

NA62

CEDAR maintenance needed -> [GASSYSTEMS-609] CEDAR maintenance needed - CERN Central Jira

change of pressure sensor used for leak warning? -> [GASSYSTEMS-741] NA62-CEDAR move to keller pressure sensor? - CERN Central Jira

 

GIF++ TGC gas system and return flow: [GASSYSTEMS-624] GIF++ TGC gas system return flow - CERN Central Jira

[GASSYSTEMS-674] MFCs to be checked - CERN Central Jira :

 

non-LHC safety valve --> check lista di Fabio

[GASSYSTEMS-172] non-LHC safety valves to be verified - CERN Central Jira

ELOG - safety valves to be verified for non-LHC systems

 

 

 

Additional requests/projects/discussions

887 CEDAR new "mixers" --> [GASSYSTEMS-509] CEDAR for beam lines - CERN Central Jira

 

AMBER

EDMS leak test? 

test certificate x-ray?

ATEX certificate?

who is technical coordinator

resource manager Oleg D?

no money? should we stop? and resume in 2026? 

Commissioning when? -> TICKET 

when they would like to start? Moritz? -> ask

 

LEAK DETECTORS LIST -> TICKET

MAINTENANCE INTERVENTIONS WORK ORDER FOR STATISTICS -> TICKET

 

PROJECT 2025

summary TICKET

 

 

examples of topics for the common discussions:

 

-cms id flushing -> Andrea

-cms r134a operation and spare -> Andrea

-cms cf4 operation and spare -> Andrea

-atl tgc nc5h12 -> Kacper

-Analysis boxes -> LPh

-C4F10 new recuperation installation -> Patrick

 

There are minutes attached to this event. Show them.
    • 09:15 → 09:20
      Piquet interventions 5m
      Speakers: Gianluca Rigoletti (CERN), persons on piquet
    • 09:20 → 09:25
      piquet planning for xmas 2026 holidays 5m

      xmas 2026 holidays: 23/12-5/1 included
      who is available for piquet?
      Gianluca 23-25
      Davide probably home, but to check
      Louis-Philippe ok
      Andrea ok -> not in the region 26-27/12
      Maria Cristina no
      Patrick ok
      Beatrice: if possible only not the last 2-3 days of holidays

      Speaker: Roberto Guida (CERN)

      MAC ok

      Gianluca ok

      Beatrice 25+or 26-> 28

      Andrea not ok (not in the region 26 and 27) -> 28->31st

      LPh ok

      Patrick ok

    • 09:25 → 09:35
      gas systems status 10m
    • 09:35 → 10:05
      news for the team
      • 09:35
        ALICE news for the team 5m
        Speakers: Louis-Philippe De Menezes (CERN), Miguel Goncalo Alves Ramos Da Silva Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT))
      • 09:40
        ATLAS news for the team 5m
        Speakers: Davide Casarotto (CERN), Maximilian Lukas Kerker (Rheinisch Westfaelische Tech. Hoch. (DE))

        ATLAS-RPC request standard RPC-CO2 from 10/5 to 16/10. to remind them that min 5 days are already needed to get good mixture composition

        ATLAS RPC does not yet want to switch to STD mixture

        Wednesday/Thursday: Eletrical work on RPC 24 V line by Patrick Villeton. Should be transparent to us.

      • 09:45
        CMS news for the team 5m
        Speakers: Andrea D'Auria (CERN), Maria Cristina Arena (CERN)
      • 09:50
        LHCb news for the team 5m
        Speakers: Patrick Carrie (CERN), Stefania Juks (Université Paris-Saclay (FR))
      • 09:55
        CLOUD news for the team 10m
        Speaker: Louis-Philippe De Menezes (CERN)

        Roberto's Notes based on discussions on October 1st

        • Software:  

          • PLC triggered early, while WinCC OA was delayed/late.

          • Updates / changes related to software to be done asap.
          • MIST soft protection
          • Server and connection issues encountered.
          • -> software should be fixed today. test with louis-philippe will follow

         

        • Repairs & Technical Issues:

          • 1 broken membrane repaired; water went inside the chamber during clean up run.

          • -MIST added electrical protection
          • - modification of transfer line today Fabio Louis-philippe today

         

        • Availability & Next Steps:

          • Marc / Gianluca available next week.

          • Follow-up scheduled for Monday.

         

        -test run expected for this weekend to check all good

    • 10:05 → 10:15
      Review new ticket or updated during current week 10m

      Jira dashboard: https://its.cern.ch/jira/secure/RapidBoard.jspa?rapidView=9098&view=detail
      Confluence page: https://confluence.cern.ch/display/KBFGS

      Speaker: Roberto Guida (CERN)
      • Try to order a new filter
      • Consider replacing filters with filters from CMS everywhere
    • 10:15 → 10:20
      other M&O news 5m
      Speakers: Beatrice Mandelli (CERN), Gianluca Rigoletti (CERN)
    • 10:20 → 10:25
      Gas mixture analysis for LHC experiments 5m

      https://epdt-gasmonitoring.web.cern.ch/

      Speaker: Maria Cristina Arena (CERN)
    • 10:25 → 10:35
      projects ready for workshop (and to be planned) 10m
      Speakers: Albin Mekkattukulam Joshy (University of Applied Sciences Bremerhaven (AF)), Amin Bouzaiene (EPFL - Ecole Polytechnique Federale Lausanne (CH)), Andrea D'Auria (CERN), Beatrice Mandelli (CERN), Davide Casarotto (CERN), Fabio Bordon, Gianluca Rigoletti (CERN), Louis-Philippe De Menezes (CERN), Maria Cristina Arena (CERN), Maximilian Lukas Kerker (Rheinisch Westfaelische Tech. Hoch. (DE)), Miguel Goncalo Alves Ramos Da Silva Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT)), Patrick Carrie (CERN), Roberto Guida (CERN), Stefania Juks (Université Paris-Saclay (FR))

       

      -mixer for Florian (Gianluca): on going

      -NA62 straw chambers disconnection (between today and friday 25->9 with fabio) not possible before monday october 12

      -ATLAS: TGC Pump 1 needs to be maintained after we switched the pump

    • 10:35 → 10:40
      FSU planning for coming weeks & Jobs status 5m
      Speaker: Andrea D'Auria (CERN)
    • 10:40 → 11:05
      EAM/Infor: update
      Convener: Miguel Goncalo Alves Ramos Da Silva Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT))
      • 10:40
        interventions from past week 5m
        Speakers: Emiliano Vaccaro (Politecnico di Torino (IT)), Miguel Goncalo Alves Ramos Da Silva Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT))

        -interventions per type (preventive, corrective, ...)

        -interventions per module

        -intervention per component

        -part used during intervention

        -time spent on intervention and total

        -divide between working hours and piquet

         

        -develop problem code list

      • 10:45
        maintenance and spare parts follow-up 5m
        Speaker: Miguel Goncalo Alves Ramos Da Silva Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT))
      • 10:50
        pump working hours 5m
        Speaker: Miguel Goncalo Alves Ramos Da Silva Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT))
    • 11:05 → 11:10
      LS3 planning preparation 5m
    • 11:10 → 11:20
      aob and requests for jobs at workshop 10m
      Speakers: Albin Mekkattukulam Joshy (University of Applied Sciences Bremerhaven (AF)), Amin Bouzaiene (EPFL - Ecole Polytechnique Federale Lausanne (CH)), Andrea D'Auria (CERN), Andrea Silvan (Universita degli Studi di Trento and INFN (IT)), Beatrice Mandelli (CERN), Bjorn Pascal Wagner, Clea Fourreaux (Université de Savoie Mont Blanc (FR)), Davide Casarotto (CERN), Emiliano Vaccaro (Politecnico di Torino (IT)), Fabio Bordon, Gianluca Rigoletti (CERN), Henrietta Szalai (Ludwig Maximilians Universitat (DE)), Louis-Philippe De Menezes (CERN), Maria Cristina Arena (CERN), Matteo Capuzzo, Maximilian Lukas Kerker (Rheinisch Westfaelische Tech. Hoch. (DE)), Michael Vogl (Fachhochschule Regensburg (DE)), Miguel Goncalo Alves Ramos Da Silva Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT)), Olaoluwa Emmanuel Olaofe (Aston University (GB)), Patrick Carrie (CERN), Roberto Guida (CERN), Ryan Oatley (University of Leeds (GB)), Stefania Juks (Université Paris-Saclay (FR)), Thomas Kattikanayil Thomas (Technical University of Liberec (CZ))
    • 11:20 → 11:25
      Matteo Capuzzo 5m
      Speaker: Matteo Capuzzo